# Recursive Language Models

### 31 Dec 2025 v1

@@ 3 Results and Discussion
Observation 1: RLMs can scale to the 10M+ token regime and can outperform base LMs and existing task-agnostic agent scaffolds on long context tasks. Across all tasks, RLMs demonstrate strong performance on input tasks well beyond the effective context window of a frontier LM, outperforming base models and common long-context scaffolds by up to $2\times$ the performance while maintaining comparable or cheaper average token costs. Notably, RLMs scale well to the theoretical costs of extending a base model’s context window – on BrowseComp-Plus (1K), the cost of GPT-5-mini ingesting 6-11M input tokens is $\mathdollar 1.50-\mathdollar 2.75$, while RLM(GPT-5) has an average cost of $\mathdollar 0.99$ and outperforms both the summarization and retrieval baselines by over $29\%$.
@@
Furthermore, on tasks where processing costs scale with the input context, RLMs make significant improvements over the base model on tasks that fit well within the model’s context window. On OOLONG, the RLM with GPT-5 and Qwen3-Coder outperform the base model by $28.4\%$ and $33.3\%$ respectively. On OOLONG-Pairs, both GPT-5 and Qwen3-Coder make little progress with F1 scores of $0.1\%$, while the RLM using these models achieve F1 scores of $58.00\%$ and $23.11\%$ respectively, highlighting the emergent capability of RLMs to handle extremely information-dense tasks.
@@
Observation 2: The REPL environment is necessary for handling long inputs, while the recursive sub-calling of RLMs provides strong benefits on information-dense inputs. A key characteristic of RLMs is offloading the context as a variable in an environment $\mathcal{E}$ that the model can interact with. Even without sub-calling capabilities, our ablation of the RLM is able to scale beyond the context limit of the model, and outperform the base model and other task-agnostic baselines on most long context settings. On the CodeQA and BrowseComp+ tasks with Qwen3-Coder, this ablation is able to outperform the RLM by $17.9\%$ and $3\%$ respectively.
@@
Observation 3: LM performance degrades as a function of input length and problem complexity, while RLM performance scales better. The benchmarks S-NIAH, OOLONG, and OOLONG-Pairs contain a fixed number of tasks over a context with lengths ranging from $2^{13}$ to $2^{18}$. Furthermore, each benchmark can be loosely categorized by different processing costs of the input context with respect to length (roughly constant, linear, and quadratic respectively). In Figure 1, we directly compare an RLM using GPT-5 to base GPT-5 on each task – we find that GPT-5 performance degrades significantly faster for more complex tasks, while RLM performance degrades but at a much slower rate, which aligns with the findings of  Goldman et al. (2025). For context lengths beyond $2^{14}$, the RLM consistently outperforms GPT-5.
@@
Furthermore, RLM costs scale proportionally to the the complexity of the task, while still remaining in the same order of magnitude of cost as GPT-5 (see Figure 9 in Appendix C). In §3.1, we explore what choices the RLM makes in these settings that causes these differences in cost. Lastly, in this setting, we also observe that the base LM outperforms RLM in the small input context regime. By construction, an RLM has strictly more representation capacity than an LM: the choice of an environment that calls the root LM is equivalent to the base LM; in practice, however, we observe that RLM performance is slightly worse on smaller input lengths, suggesting a tradeoff point between when to use a base LM and when to use an RLM.
@@
Observation 4: The inference cost of RLMs remain comparable to a base model call but are high variance due to differences in trajectory lengths. RLMs iteratively interact with their context until they find a suitable answer, leading to large differences in iteration length depending on task complexity. In Figure 3, we plot the quartile costs for each method across all experiments in Table 1 excluding BrowseComp-Plus (1K), as the base models cannot fit any of these tasks in context. For GPT-5, the median RLM run is cheaper than the median base model run, but many outlier RLM runs are significantly more expensive than any base model query. However, compared to the summarization baseline which ingests the entire input context, RLMs are up to $3\times$ cheaper while maintaining stronger performance across all tasks because the model is able to selectively view context.
@@
We additionally report runtime numbers of each method in Figures 56 in Appendix C, but we note several important caveats. Unlike API costs, these numbers are heavily dependent on implementation details such as the machine used, API request latency, and the asynchrony of LM calls. In our implementation of the baselines and RLMs, all LM calls are blocking / sequential. Nevertheless, similar to costs, we observe a wide range of runtimes, especially for RLMs.
@@
Observation 5: RLMs are a model-agnostic inference strategy, but different models exhibit different overall decisions on context management and sub-calling. While GPT-5 and Qwen3-Coder-480B both exhibit strong performance as RLMs relative to their base model and other baselines, they also exhibit different performance and behavior across all tasks. On BrowseComp-Plus in particular, RLM(GPT-5) nearly solves all tasks while RLM(Qwen3-Coder) struggles to solve half.
@@
We note that the RLM system prompt is fixed for each model across all experiments and is not tuned for any particular benchmark. Between GPT-5 and Qwen3-Coder, the only difference in the prompt is an extra line in the RLM(Qwen3-Coder) prompt warning against using too many sub-calls (see Appendix D). We provide an explicit example of this difference in example B.3, where RLM(Qwen3-Coder) performs the semantic transformation in OOLONG as a separate sub-LM call per line while GPT-5 is conservative about sub-querying LMs.
@@ 3.1 Emergent Patterns in RLM Trajectories
Even without explicit training, RLMs exhibit interesting context management and problem decomposition behavior. We select several examples of snippets from RLM trajectories to understand how they solve long context problems and where they can improve. We discuss particular examples of interesting behavior here, with additional examples in Appendix B.
Figure 4: RLMs have common patterns in their trajectories when solving tasks. (a) We frequently observed RLMs filtering and interacting with their context through code like regex queries. (b) We found that RLMs can effectively decompose their context through recursive sub-calls (c) On long-output tasks, RLMs are able to solve sub-problems using recursive sub-LM calls and stitch their outputs to form a final output.
Filtering input information using code execution based on model priors. A key intuition for why the RLM abstraction can maintain strong performance on huge inputs without exploding costs is the LM’s ability to filter input context without explicitly seeing it. Furthermore, model priors enable the RLM to narrow the search space and process fewer input tokens. As an example, in Figure 4a, we observed RLM(GPT-5) using regex queries search for chunks containing keywords in the original prompt (e.g. “festival”) and phrases it has a prior about (e.g. “La Union”). Across most trajectories, a common strategy we observed was probing the context by printing a few lines back to the root LM, then filtering based on its observations.
Chunking and recursively sub-calling LMs. RLMs defer essentially unbounded-length reasoning chains to sub-(R)LM calls. The choice of decomposition can greatly affect task performance, especially for information-dense problems. In our experiments, we did not observe complicated partitioning strategies beyond uniform chunking or keyword searches. In Figure 4b, RLM(Qwen3-Coder) chunks by newline in a 1000+ line context from OOLONG.
Answer verification through sub-LM calls with small contexts. We observed several instances of answer verification made by RLMs through sub-LM calls. Some of these strategies implicitly avoid context rot by using sub-LMs to perform verification (see example B.1), while others solely use code execution to programmatically verify answers are correct. In some instances, however, the answer verification is redundant and significantly increases the cost per task — in example B.3, we observed a trajectory on OOLONG where the model tries to reproduce its correct answer more than five times before choosing the incorrect answer in the end.
@@ 4 Related Works
Long Context LM Systems. There have primarily been two orthogonal directions for long context management in language model systems: 1) directly changing the architecture of and retraining the base LM to handle longer contexts (Press et al., 2022; Gu et al., 2022; Munkhdalai et al., 2024), and 2) building a scaffold around the LM that implicitly handles the context – RLMs focus on the latter. One popular class of such strategies is lossy context management, which uses summarization or truncation to compress the input context at the cost of potentially losing fine-grained information. For example, MemWalker (Chen et al., 2023) constructs a tree-like data structure of the input that the LM can navigate when answering long context questions. ReSum (Wu et al., 2025) is another work that adds a summarization tool to periodically compress the context of a multi-turn agent. Another class of strategies implement an explicit memory hierarchy in the agent scaffold (Packer et al., 2024; Chhikara et al., 2025; Zhang et al., 2025). RLMs are different from prior work in that all context window management is implicitly handled by the LM itself.
Task Decomposition through sub-LM calls. Many LM-based agents (Guo et al., 2024; Anthropic, 2025) use multiple, well-placed LM calls to solve a problem, however many of these calls are placed based on human-engineered workflows. Several methods like ViperGPT Surís et al. (2023), THREAD (Schroeder et al., 2025), DisCIPL (Grand et al., 2025), ReDel Zhu et al. (2024), Context Folding (Sun et al., 2025), and AgentFold (Ye et al., 2025) have explored deferring the choice of sub-LM calls to the LM. These techniques emphasize task decomposition through recursive LM calls, but are unable to handle long context inputs beyond the length of the base LM. RLMs, on the other hand, are enabled by an extremely simple intuition (i.e., placing the prompt as part of the external environment) to symbolically manipulate arbitrarily long strings and to iteratively refine their recursion via execution feedback from the persistent REPL environment.
@@ 5 Limitations and Future Work
While RLMs show strong performance on tasks beyond the context window limitations of existing LMs at reasonable inference costs, the optimal mechanism for implementing RLMs remains under-explored. We focused on synchronous sub-calls inside of a Python REPL environment, but we note that alternative strategies involving asynchronous sub-calls and sandboxed REPLs can potentially significantly reduce the runtime and inference cost of RLMs. Furthermore, we chose to use a max recursion depth of one (i.e. sub-calls are LMs); while we found strong performance on existing long-context benchmarks, we believe that future work should investigate deeper layers of recursion.
Lastly, we focused our experiments on evaluating RLMs using existing frontier models. Explicitly training models to be used as RLMs (e.g. as root or sub-LMs) could provide additional performance improvements – as we found in §3.1, current models are inefficient decision makers over their context. We hypothesize that RLM trajectories can be viewed as a form of reasoning (OpenAI et al., 2024; DeepSeek-AI et al., 2025), which can be trained by bootstrapping existing frontier models (Zelikman et al., 2022; 2024).
@@ 6 Conclusion
We introduced Recursive Language Models (RLMs), a general inference framework for language models that offloads the input context and enables language models to recursively sub-query language models before providing an output. We explored an instantiation of this framework that offloads the context into a Python REPL environment as a variable in memory, enabling the LM to reason over its context in code and recursive LM calls, rather than purely in token space. Our results across multiple settings and models demonstrated that RLMs are an effective task-agnostic paradigm for both long-context problems and general reasoning. We are excited to see future work that explicitly trains models to reason as RLMs, which could result in another axis of scale for the next generation of language model systems.
@@ Acknowledgments
This research is partially supported by the Laude Institute. We thank Noah Ziems, Jacob Li, James Moore, and the MIT OASYS and MIT DSG labs for insightful discussions throughout this project. We also thank Matej Sirovatka, Ofir Press, Sebastian Müller, Simon Guo, and Zed Li for helpful feedback.
Appendix A Negative Results: Things we Tried that Did Not Work.
@@
Depending on the model, distinguishing between a final answer and a thought is brittle for RLMs. The current strategy for distinguishing between a “next turn” and a final answer for the RLM is to have it wrap its answer in FINAL() or FINAL_VAR() tags. Similar to intuition about structured outputs degrading performance, we also found the model to make strange decisions (e.g. it outputs its plan as a final answer). We added minor safeguards, but we also believe this issue should be avoided altogether in the future when models are trained as RLMs.
Appendix B Additional RLM Trajectories
In this section, we provide several example trajectories to highlight characteristics of frontier models as RLMs. Many of the trajectories are too long to fit in text (we also provide the raw trajectories and a visualizer in our codebase), so we describe each step and show specific examples when relevant.
A few noticeable properties of these trajectories are that RLMs often make non-optimal choices despite their strong results in §2. For example, in Example B.2, we observed that the RLM with Qwen3-Coder carefully constructs its final answer through a mix of recursive sub-calls and code execution in the first iteration, but then discards this information and continues wasting sub-calls before not using these stored answers. We also observed distinct differences in model behavior such as in Example B.3, where we found Qwen3-Coder make hundreds to thousands of recursive sub-calls for a single simple task, while GPT-5 makes on the order of ten. While these examples are not comprehensive, they provide useful qualitative insight into how to improve RLMs.
B.1 RLM(GPT-5) on BrowseComp-Plus-Query_74
B.2 RLM(Qwen3-Coder) on OOLONG-Pairs-Query_3
B.3 RLM(Qwen3-Coder) on OOLONG-Query_212
B.4 RLM(GPT-5) on CodeQA-Query_44
Final. The RLM answers choice ‘1’, which is the correct answer.
Appendix C Additional Runtime and Cost Analysis of RLMs
We supplement the cost and runtime analysis of RLMs with additional, fine-grained plots. In Figures 7, 8 we include a histogram for the cost of each method on every task for both GPT-5 and Qwen3-Coder. We generally observe long-tailed, high-variance trajectories for RLMs in both models.
Figure 5: Plotted quartiles of the runtime GPT-5 across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §$2.$ We plot the 25th, 50th, 75th, and 95th percentiles.
Figure 6: Plotted quartiles of the runtime Qwen3-Coder-480B across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §$2.$ We plot the 25th, 50th, 75th, and 95th percentiles.
Figure 7: Histogram of the API costs for GPT-5 across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §$2.$
Figure 8: Histogram of the API costs for Qwen3-Coder-480B across OOLONG, OOLONG-Pairs, CodeQA, and BrowseComp+ (1K) for all methods described in §$2.$
Figure 9: We plot the API cost in USD for the runs in Figure 1.
Appendix D Additional Methods and Baseline Details
D.1 Prompts for Experiments
We focus on methods that are entirely task agnostic, so we fix our prompt for each method across all tasks. For the RLM prompt, the only difference between GPT-5 and Qwen3-Coder is an added line in the beginning that warns Qwen3-Coder not to use too many sub-LM calls – we found in practice that without this warning, the model will try to perform a subcall on everything, leading to thousands of LM subcalls for basic tasks! In this section, we provide the system prompt used for all methods in §2.1 (other than the base model, which does not include a system prompt).

### 28 Jan 2026 v2

@@ 4 Results and Discussion
Observation 1: RLMs can scale to the 10M+ token regime and can outperform base LMs and existing task-agnostic agent scaffolds on long context tasks. Across all tasks, RLMs demonstrate strong performance on prompts well beyond the effective context window of a frontier LM, outperforming base models and common long-context scaffolds by up to $2\times$ the performance while maintaining comparable or cheaper average token costs. Notably, RLMs scale well beyond the base models’ context window. For instance, on BrowseComp-Plus (1K), a linearly extrapolated cost for GPT-5-mini ingesting 6-11M input tokens is $\mathdollar 1.50-\mathdollar 2.75$, while RLM(GPT-5) has an average cost of $\mathdollar 0.99$ and outperforms both the summarization and retrieval baselines by over $29\%$.
@@
Furthermore, on tasks where processing costs scale with the input context, RLMs make significant improvements over the base model, even on tasks within the model’s context window. On OOLONG, the RLM with GPT-5 and Qwen3-Coder outperform the base model by $28.4\%$ and $33.3\%$ respectively. On OOLONG-Pairs, both GPT-5 and Qwen3-Coder make little progress with F1 scores of $0.1\%$, while the RLM using these models achieve F1 scores of $58.0\%$ and $23.1\%$ respectively, highlighting the emergent capability of RLMs to handle extremely information-dense tasks.
@@
Observation 2: The REPL is necessary for handling long inputs, while the recursive sub-calling of RLMs provides strong benefits on information-dense inputs. A key characteristic of RLMs is offloading the context as a variable in an environment $\mathcal{E}$ that the model can interact with. Even without sub-calling capabilities, our ablation of the RLM is able to scale beyond the context limit of the model and outperform other task-agnostic baselines on most long context settings. On the CodeQA and BrowseComp+ tasks with Qwen3-Coder, this ablation is able to outperform the RLM by $17.9\%$ and $3\%$ respectively.
@@
Observation 3: LM performance degrades as a function of input length and problem complexity, while RLM performance scales better. The benchmarks S-NIAH, OOLONG, and OOLONG-Pairs contain a fixed number of tasks over contexts with lengths ranging from $2^{13}$ to $2^{18}$. Each benchmark can be loosely categorized by different processing complexity of the input context with respect to length (roughly constant, linear, and quadratic respectively). In Figure 1, we directly compare an RLM using GPT-5 to base GPT-5 on each task. We find that GPT-5 performance degrades significantly faster for more complex tasks, while RLM performance degrades at a much slower rate, which aligns with the findings of Goldman et al. (2025). For context lengths beyond $2^{14}$, the RLM consistently outperforms GPT-5.
@@
Furthermore, RLM costs scale proportionally to the complexity of the task, while still remaining in the same order of magnitude of cost as GPT-5 (see Figure 11 in Appendix F). In §4.1, we explore the choices that the RLM makes that cause these differences in cost. Lastly, in this setting, we also observe that the base LM outperforms RLM in the small input context regime. By construction, a RLM has strictly more representation capacity than an LM. In practice, however, we observe that RLM performance is slightly worse on smaller input lengths, suggesting a tradeoff point between when to use a base LM and when to use an RLM.
@@
Observation 4: The inference cost of RLMs remains comparable to a base LM call but has high variance due to differences in trajectory lengths. RLMs iteratively interact with their context until they find a suitable answer, leading to large differences in iteration length depending on task complexity. In Figure 3, we plot the quartile costs for each method across all experiments in Table 1 excluding BrowseComp-Plus (1K), as the base models cannot fit any of these tasks in context. For GPT-5, the median RLM run is cheaper than the median base model run, but many outlier RLM runs are significantly more expensive than any base model query. However, compared to the summarization agent which ingests the entire input context, RLMs are up to $3\times$ cheaper while maintaining stronger performance across all tasks because the RLM is able to selectively view context.
@@
We additionally report runtime numbers of each method in Figures 78 in Appendix F, but we note several important caveats. Unlike API costs, these numbers are heavily dependent on implementation details such as the machine used, API request latency, and the asynchrony of LM calls. In our implementation of the baselines and RLMs, all LM calls are blocking / sequential. Nevertheless, similar to costs, we observe a wide range of runtimes, especially for RLMs.
@@
Observation 5: RLMs are a model-agnostic inference strategy, but different models exhibit different overall decisions on context management and sub-calling. While GPT-5 and Qwen3-Coder-480B both exhibit strong performance as RLMs relative to their base model and other baselines, they also exhibit different performance and behavior across all tasks. On BrowseComp-Plus (1k) in particular, RLM(GPT-5) nearly solves all tasks while RLM(Qwen3-Coder) struggles to solve half.
@@
We note that the RLM system prompt is fixed for each model across all experiments and is not tuned for any particular benchmark. Between GPT-5 and Qwen3-Coder, the only difference in the prompt is an extra line in the RLM(Qwen3-Coder) prompt warning against using too many sub-calls (see Appendix C). We provide an explicit example of this difference in example E.3, where RLM(Qwen3-Coder) launches a sub-call per line in OOLONG while GPT-5 is conservative about sub-querying LMs.
Observation 6: Training RLMs on one domain can improve general downstream RLM performance. Certain behavior in RLM trajectories are common among different domains, such as probing the input and recursively sub-calling on shorter contexts. In Table 1, we find that RLM-Qwen3-8B, a Qwen3-8B model that we fine-tuned on RLM(Qwen3-Coder-480B-A35B) trajectories on a small, unrelated set of tasks (LongBenchPro; Chen et al. 2026) considerably outperforms the base Qwen3-8B as an RLM by $28.3\%$ on average. Furthermore, its inference costs are much lower due to better decision making and fewer mistakes as an RLM.
@@ 4.1 Emergent Patterns in RLM Trajectories
Even without explicit training, RLMs exhibit interesting context and problem decomposition behavior. We select several examples of snippets from RLM trajectories to understand how they solve long context problems and where they can improve. We discuss particular examples of interesting behavior here, with additional examples in Appendix E.
Figure 4: RLMs have common patterns in their trajectories when solving tasks. (a) We frequently observed RLMs filtering and interacting with their context through regex code. (b) We found that RLMs can effectively decompose their context through recursive sub-calls (c) On long-output tasks, RLMs are able to solve sub-problems using recursive sub-LM calls and stitch their outputs to form a final output.
Chunking and recursively sub-calling LMs. RLMs defer essentially unbounded-length reasoning chains to sub-LM calls. The choice of decomposition can greatly affect task performance, especially for information-dense problems. In our experiments, we did not observe complicated partitioning strategies beyond uniform chunking or keyword searches. In Figure 4b, RLM(Qwen3-Coder) chunks by newline in a 1000+ line context from OOLONG.
Filtering input information using code execution based on model priors. A key intuition for why the RLM abstraction can maintain strong performance on huge inputs without exploding costs is the LM’s ability to filter input context without explicitly seeing it. Furthermore, model priors enable the RLM to narrow the search space and process fewer input tokens. As an example, in Figure 4a, we observed RLM(GPT-5) using regex queries to search for chunks containing keywords in the original prompt (e.g. “festival”) and phrases it has a prior about (e.g. “La Union”).
@@ 5 Related Works
Long-Context LM Systems. There have primarily been two orthogonal directions for long-context management in language model systems: 1) directly changing the architecture of and retraining the base LM to handle longer contexts (Press et al., 2022; Gu et al., 2022; Munkhdalai et al., 2024), and 2) building a scaffold around the LM that implicitly handles the context – RLMs focus on the latter. One popular class of such strategies is lossy context management, which uses summarization or truncation to compress the input context at the cost of potentially losing fine-grained information. For example, MemWalker (Chen et al., 2023) constructs a tree-like data structure of the input that the LM can navigate when answering long context questions. ReSum (Wu et al., 2025) is another work that adds a summarization tool to periodically compress the context of a multi-turn agent. Another class of strategies implement an explicit memory hierarchy in the agent scaffold (Packer et al., 2024; Chhikara et al., 2025; Zhang et al., 2025). RLMs differ from these works in that all context window management is implicitly handled by the LM itself.
Task Decomposition through sub-LM calls. Many LM-based agents (Guo et al., 2024; Anthropic, 2025) use multiple, well-placed LM calls to solve a problem; however, many of these calls are placed based on human-engineered workflows. Several methods like ViperGPT (Surís et al., 2023), THREAD (Schroeder et al., 2025), DisCIPL (Grand et al., 2025), ReDel (Zhu et al., 2024), Context Folding (Sun et al., 2025), and AgentFold (Ye et al., 2025) have explored deferring the choice of sub-LM calls to the LM. These techniques emphasize task decomposition through recursive LM calls, but are unable to handle long context inputs beyond the length of the base LM. RLMs, on the other hand, are enabled by an extremely simple intuition (i.e., placing the prompt in the external environment) to symbolically manipulate arbitrarily long strings and to iteratively refine their recursion via execution feedback from the persistent REPL.
@@ 6 Limitations and Future Work
While RLMs show strong performance on tasks beyond the context window limitations of existing LMs at reasonable inference costs, evaluations for more difficult and natural long-context processing tasks and the best mechanisms for implementing RLMs both remain highly under-explored. We focused on synchronous sub-calls inside of a Python REPL environment, but we note that alternative strategies involving asynchronous sub-calls and sandboxed REPLs can potentially significantly reduce the runtime and inference cost of RLMs. Furthermore, we chose to use a max recursion depth of one (i.e. sub-calls are LMs); while we found strong performance on existing long-context benchmarks, we believe that future work should investigate deeper levels of recursion or even new hybrids between symbolic recursion and neural attention. We include additional limitations and negative results in Appendix B.
Lastly, we focused our experiments on evaluating RLMs using existing frontier models, but show initial evidence on a Qwen3-8B model that explicitly training a model to be used as a RLM provides very rapid performance improvements, even outside the training domain. We hypothesize that RLM trajectories can be viewed as a form of reasoning (OpenAI et al., 2024; DeepSeek-AI et al., 2025), which can be trained by bootstrapping existing models (Zelikman et al., 2022, 2024). We hope that training native RLMs can be treated as a new axis of scale to improve LM performance on general and long-horizon tasks.
@@ 7 Conclusion
We introduced Recursive Language Models (RLMs), a general inference framework for language models that offloads the input context and enables language models to recursively sub-query language models before providing an output. We explored an instantiation of this framework that offloads the context into a Python REPL environment as a variable in memory, enabling the LM to reason over its context in code and recursive LM calls, rather than purely in token space. Our results across multiple settings and models demonstrated that RLMs are an effective task-agnostic paradigm for both long-context problems and general reasoning. Building on our small fine-tuning experiments, we are excited to see future work that explicitly trains models to reason as RLMs, which could result in another axis of scale for the next generation of language model systems.
8 Impact Statement
This paper explores a strategy for enabling language models to solve long context problems and scaling future language model systems. The goal is to advance research on systems that can help us solve complex problems. While there are potential societal consequences of this work, we believe they are not specific to this paper and do not need to be highlighted here.
@@ Acknowledgments
This research is partially supported by the Laude Institute, Prime Intellect, and Modal Labs. We thank Noah Ziems, Jacob Li, James Moore, and the MIT OASYS and MIT DSG labs for insightful discussions throughout this project. We also thank Jack Cook, Matej Sirovatka, Ofir Press, Sebastian Müller, Simon Guo, and Zed Li for helpful feedback.
Appendix A Additional Training Details
We trained RLM-Qwen3-8B as a very small scale exercise in training the first natively recursive language model. We hypothesized that, though acting as an RLM appears to produce sophisticated behavior due to recursion, it can be sufficient to focus on improving the root LM’s ability to interact with the programmatic representation of the prompt in the REPL and to discern when sub-calls are useful. In other words, while a typical RLM trajectory can be extremely long due to all of the sub-calls potentially launched (possibly $\Omega(|P|)$ for a prompt $P$), the leaf sub-calls are essentially general-purpose LLM requests and the major hurdle is learning to operate as the root model.
This simple insight allowed us to explore a similarly simple recipe for training. In particular, we sampled RLM trajectories from a larger language model (Qwen3-Coder-480B-A35B-Instruct; Qwen Team 2025b) and, after filtering, distilled them to a smaller model (Qwen3-8B; Qwen Team 2025a) from the same model family. We evaluated RLM(Qwen3-Coder-480B-A35B) on 750 English LongBenchPro (Chen et al., 2026) tasks, collecting a total of 2250 candidate trajectories.
We first remove trajectories that score exactly 0.0 on the benchmark or do not go beyond one turn, bringing it down to 1,072 candidate trajectories. We separated each root RLM turn (i.e. iteration) as a separate SFT sample consisting of an input (the full history) and output (the output the root LM gave at that step).
We then applied a filtering step to remove turns beyond the context limit of Qwen3-8B (we approximated this as 100k characters), and also applied an extra programmatic correction step to fix small template mistakes in RLMusage (e.g. outputting final answers, calling the REPL, etc.). To elaborate, we noticed that trajectories generated by Qwen3-Coder-480B-A35B had noticeable mistakes in following the RLM instructions, which hurt the performance of the distilled RLM-Qwen3-8B. For example, it would often mix FINAL(answer) with FINAL(variable in REPL). We added an extra programmatic fixing step to look for common templated mistakes and patch them, leading to much better performance in the final RLM-Qwen3-8B. In total, 16% of turns cleaned incorrectly used FINAL answers, and 13% of turns incorrectly called a variable from the REPL (i.e. FINAL_VAR) as a final answer. In Figure 5, we show pre- and post-filtering statistics for our training trajectories.
Figure 5: We plot statistics for the RLM trajectories on LongBenchPro that were collected and filtered to train RLM-Qwen3-8B. The left plots show the unfiltered trajectories, and right plots show the post-filtering trajectories.
We used the prime-rl library (Intellect, 2025) for fine-tuning. We used a batch size of 64 for 300 training steps, training for 48 H100 hours. While this exceedingly simple training recipe was able to demonstrate substantial gains for our 8B model, we call on future work to investigate training native RLMs much more thoroughly. We expect that doing so at much larger scales in terms of model size, number and variety of examples, and number of (ideally on-policy and online) rollouts will be necessary to maximize the potential of RLMs.
Appendix B Negative Results: Things we Tried that Did Not Work.
@@ Appendix A Additional Training Details
Depending on the model, distinguishing between a final answer and a thought is brittle for RLMs. The current strategy for distinguishing between a “next turn" and a final answer for the RLM is to have it wrap its answer in FINAL() or FINAL_VAR() tags. Similar to intuition about structured outputs degrading performance, we also found the model to make strange decisions (e.g. it outputs its plan as a final answer). We added minor safeguards, but we also believe this issue should be avoided altogether in the future when models are trained as RLMs.
Appendix C Additional Methods and Baseline Details
C.1 Prompts for Experiments
We focus on methods that are entirely task agnostic, so we fix our prompt for each method across all tasks. For the RLM prompt, the only difference between GPT-5 and Qwen3-Coder is an added line in the beginning that warns Qwen3-Coder not to use too many sub-LM calls – we found in practice that without this warning, the model will try to perform a subcall on everything, leading to thousands of LM subcalls for basic tasks! For the fine-tuned Qwen3-8B experiment, we provide a slightly different prompt due to the differences in context window size of the smaller model (from 272k to 32k). In this section, we provide the system prompt used for all methods in §3.1 (other than the base model, which does not include a system prompt).